Two consecutive integers add to make 315.
Jamal writes expressions for the two integers:
1st number n, 2nd number n+2
a) Add Jamal's expressions together.
b) he two integers add to make 315. Write this as an equation using your answer to part a.
c) Solve your equation to find the value of a.
d) Write down the two integers.

Answer by math_helper(2461)

Two consecutive integers add to make 315.
Jamal writes expressions for the two integers:
1st number n, 2nd number n+2
------------------
"Two consecutive integers" should be n and n+1 (not n and n+2)


a) n + (n+1) = 2n+1
b) 2n + 1 = 315
c) 2n + 1 = 315
Subtract 1 from both sides, then...
Divide both sides by 2...
now you will have the value of n
Step (d) is then trivial